FEL Upgrade Project Weekly Brief
March 5-9, 2001


For the last week of the winter FEL user run , the FEL produced new power records for femtosecond class green light (56 watts) and blue (350nm) light (16 watts). This light was used for patterning experiments on photosensitive glass by our user from the Aerospace Corp.   We also obtained very nice angular distributions of the Compton x-ray source.

Please check out the new and  improved web site http://laser.jlab.org there are images and movies from the last run in the "media gallery" located on the top of the index page. Thanks to Kevin Jordan and Michelle Shinn for this production.


A draft rebaseline of the Phase 1 cost performance plan and the initial plan for the Phase 2 project was forwarded to the ONR contract monitor for review.   The Phase 1 was rebaselined to reflect  Engineering Change Requests approved at the October 2000 project review:  (1) the descoping of the optical cavity from an R5 configuration to a simpler and less costly concentric cavity; (2) the evolution from the driver accelerator
design from Release 1.0 to Release 1.1; and the (3) the change in vacuum hardware from 2 inch to 3 inch aperture. The initial Phase 2 cost performance plan has been been drafted to optimize the schedule and
minimize the cost of the complete Phase 1 and Phase 2 efforts.

On Friday March 9 we had a visit from a team from AES who presented a summary of their SBIR effort funded by SMDC on the design of a 100 mA injector based on the current 5 mA FEL injector.

WBS 4 (Injector):

Continued work on refining the gun vacuum chamber and cesiation scheme. Ordered flanges and ion pumps for the gun chamber.

WBS 8 (I&C):

Designs are nearly complete for the User Lab LAN Patch Panels and the Hutch Video Patch panels. The last two LSS Lab Interface Boxes are under construction and progressing nicely. One of these will be installed in User Lab #2 to secure the lab for mirror testing, and the other will go in "lab #7" for optics instrumentation development work. Progress continues to be made on the Optical Beam Position Monitors (OBPMs), time will be scheduled after the run to continue the development.

Many thanks to the software group for their work on the new trim card specification, which can be found at:
http://devweb.acc.jlab.org/controls_web/LLAPPS/Requirements/FelTrim.html. The new system will be capable of up to 20 amps per channel, each card will have hardware strapping for full scale outputs of 1, 5, 10 or 20 amps maximum current.  This system will be purchased commercially.

Parts are on order for 40 new Ion Pump Power Supplies. The crate design is also progressing. Discussions continue for a redesign of the fast valve cards, this would assist the CEBAF support group since there are no spares and the programmable devices used are long since out of date.

First Draft Design Specifications for a new VME based timing module is complete. This will be circulated and then proceed to the design phase. This card will provide timing signals for various sample & holds, and instrumentation devices on both the electron beam transport and the FEL.

Four (4) Drawings have been completed and submitted to Document Control.  EECAD has the Picomotor Power supply Chassis, User Lab Power Supply Chassis and the Beam Viewer Interlock Box for documentation.  Documentation continues on the Video Crosspoint Chassis wiring diagram, Alignment
Mode Permit Card Schematic, and the Lab Status Lamp Control Chassis.  50 VME front panels for various active projects were shipped out for machining and silk-screening this week and are expected back in about 30 days.  Another 50 will be ordered today to complete this project.

WBS 9 (Transport):

Injector Dipoles (DU/DV)
o  DULY Research have now obtained a field that was within the flatness specification using their 3D model
    (in RADIA) of the actual wedge model of the GV dipole but with thicker mu metal, indicating that their
    still-to-large mesh for the mu metal is the cause of the gradient reported earlier.  This has given them
    confidence to increase that mesh.  They are contemplating buying a new 800 MHz, large memory
    computer to handle the increased number of nodes.  Their drawing set and the drafting model of the GV
    (Small injector dipole) has now added the coil leads (pdf file enclosed).
Optical Chicane Dipole (DW)
o  We continued back check of the remaining parts, the package should be ready in two weeks.
Arc Dipoles (GY, GX, GQ)
o  AES continues progress in designing GY (180 degree dipole) coil's lead area
o  AES obtained good results with their 3D magnetic model of the GY. Some additional fine structure to the
    mesh is required to investigate the fields down to the 1 part in 10,000 level.
QX (3.125" Quad)
o Measurement Probe: The probe was glued together in its final configuration and should be ready Monday.
o Design Package: The QX drawing package has had no additional work done on it because Jeff Karn has
    been directed to work on a critical path Hall A magnet.
Sextupole/Octupole /Corrector
o  We decided that the Octupole would most likely not be able to be generated from additional coil s added to
    the sextupole because of the huge quadrupole generated.  It will most likely be a separate magnet. Robin
    Wines continued work on the magnetic model of the Sextupole, attempting to obtain a vertical corrector
    dipole with the added coil method.  This attempt at a combination magnet was also not encouraging because
    of  the large sextupole in the 2D model.  The corrector was 1 part in 100 only within half of the required
    width.  (A 3D model may reveal a compensating counter sextupole from the end fields that would extend the
    good field region).  The remaining time that Robin has before the expected birth of her baby in three weeks
    will be devoted to perfection and analysis of the sextupole.  We have decided to send design of the
    Octupole out to an Engineering Services Contractor.
o  Another round of distributing diagnostics and correction elements into the lattice in consultation with David
    Douglas, our chief  beam physicist, simplified the component packaging process further and fixed the
    number  of required correctors and their type to an extent that we can start the ordering process.
o We started work on defining the engineering services contractor task order that will generate the design of the
    dipole vacuum chambers.

WBS 10 (Wiggler):

We received the Dispersion Section for the Optical Klystron from Process Equipment Co.  It looks good.

WBS 11 (Optics):

We received the vacuum vessel for the mirror test stand, a major component for our prototype cavity mirror tests.  We will check it against the prints to confirm dimensions.  We have new concepts for installing 4 mirrors in the optical cavity assemblies, and are working with the designers to see how feasible they are.  An overall system print, defining the optics subsystem regions, and positions for components, was reviewed and signed

Optics supported operations for ourselves (harmonic conversion of 1 micron FEL light), H. Helvajian (irradiation of doped glass), J. Boyce (x-ray studies), I&C (O-BPM detector tests) and M. Kelley (evaluation of Ge for pulse switching).


The harmonic conversion results are record-breaking (for ultrafast laser output in the visible and near-UV), and therefore deserve some additional comment.  We found that the second harmonic generation (SHG) of 1.05 micron FEL light using lithium triborate (LBO) was ~ 40% and constant up to 185 W of input power.  Maximum visible output, corrected for optics, was ~ 56W.  There was no apparent damage.  To convert to the UV, two different schemes were used.  For third harmonic generation (THG) at 350 nm, the SHG and fundamental were incident on another LBO crystal.  For fourth harmonic generation (FHG) at 262.5 nm, the SHG output was focused, then directed onto a beta barium borate (BBO) crystal.  Conversion efficiencies from the IR to the UV were about the same for THG and FHG, 9% and 8% respectively. We briefly measured a THG output of 16 W for 100 W of input, but more often measured powers of 11W.  An IR camera showed that the second LBO crystal was heating up, and we observed a decrease in power.  This could be due to color centers being formed during irradiation, further study is needed to confirm this.  We did not have the opportunity to measure the FHG conversion under CW operation, but did use the pulsed  output to irradiate H. Helvajian's samples.  We also used the 350 nm CW output to irradiate several of his samples.  As was mentioned in the Dec 8 weekly report, harmonic conversion is sensitive to the input wavelength and intensity, we made some progress in monitoring both by analyzing the FEL fundamental radiation with a spectrograph and InGaAs line scan camera (essentially a linear array) and displaying this in the control room.  A lot more work needs to be done to get this into our controls system, but the components are coming together.

Despite constant distractions of descriptions of South Sea Islands from FEL Operations staff, it is becoming routine to measure Thomson backscattered X-rays from the FEL, both at ~5 keV and ~12 keV. This week the first successful diffraction pattern of these X-rays off a LiF crystal was measured.  These measurements will be used to position the diffracting crystal for maximum flux for possible future IR-Xray pump-probe experiments.

Privacy and Security Notice